# Both Direct and Indirect Evidence Contribute to Dative Alternation Preferences in Language Models

### v2

@@ Abstract
Language models (LMs) tend to show human-like preferences on a number of syntactic phenomena, but the extent to which these are attributable to direct exposure to the phenomena or more general properties of language is unclear. We explore this with the English dative alternation (DO: gave Y the X vs. PO: gave the X to Y), using a controlled rearing paradigm wherein we iteratively train small LMs on systematically manipulated input. We focus on properties that affect the choice of alternant: length and animacy. Both properties are directly present in datives but also reflect more global tendencies for shorter elements to precede longer ones and animates to precede inanimates. First, by manipulating and ablating datives for these biases in the input, we show that direct evidence of length and animacy matters, but easy-first preferences persist even without such evidence. Then, using LMs trained on systematically perturbed datasets to manipulate global length effects (re-linearizing sentences globally while preserving dependency structure), we find that dative preferences can emerge from indirect evidence. We conclude that LMs’ emergent syntactic preferences come from a mix of direct and indirect sources.
@@ 1 Introduction
Yet most people would prefer the latter. The reason is that users of English typically show a “short-first” preference in the dative alternation. That is, if the theme is longer than the recipient they prefer the DO. If the theme is shorter, the PO. See an illustration of this in Fig. 1. Besides length, there are a host of other factors that influence the choice of alternation used, such as animacy and pronominality (Bresnan et al., 2007; de Marneffe et al., 2012).
@@
Human-like preferences in the dative alternation are learned well by even relatively small LMs (Hawkins et al., 2020; Misra & Kim, 2024). But what has to be true of the training data for these preferences to emerge? Recent work has considered the role of indirect evidence in language learning, asking whether LMs learn syntactic phenomena by memorizing only the most similar instances or whether they rely on structurally similar instances in order to generalize. Evidence so far has been mixed, with some work arguing for generalization from indirect evidence. Misra & Mahowald (2024) find that a somewhat rare syntactic structure is learned, in part, from indirect evidence from more common related structures. Patil et al. (2024) find that grammatical generalizations are possible even when never encountered. Oba et al. (2024), though, find that indirect evidence is less important than direct evidence. Anonymous (2025) finds that binomial preferences (e.g., “peanut butter and jelly” vs. “jelly and peanut butter”) in models largely depends on memorization, with little evidence of generalization to human-like preferences for novel binomial pairs. Jumelet et al. (2024) find that both direct evidence and indirect evidence matter for adjective order (e.g., “big black boxes” vs. “black big boxes”).
@@
We posit that the dative alternation is an elegant testing ground for exploring these questions, since it is linguistically well-studied and shown to be sensitive to a variety of factors that are directly observable in extant dative data, but which also depend on more general constraints in the language. One possibility (direct evidence) is that LMs prefer dative constructions that are highly similar to those observed in training. That is, if She gave it to the boy who signed up for class and was excited is attested in the training but She gave the boy who signed up for class and was excited it is not, then LMs might assign higher probability to strings that are very similar to the former. A second possibility (indirect evidence) is that the observed dative preferences are emergent from more general features of the input data. Specifically, we hypothesize that the LMs’ preference for “short-first” and “animate-first” dative utterances (MacDonald, 2013) might follow not from memorizing instances of such datives but from those preferences existing more generally in the data.
@@
To adjudicate between these possibilities, we use a controlled rearing paradigm (Jumelet et al., 2021; Frank, 2023; Misra & Mahowald, 2024; Feng et al., 2024; Patil et al., 2024; Leong & Linzen, 2024; Kallini et al., 2024; Xu et al., 2025), whereby we train small models on carefully controlled input data and then test whether the predicted preferences emerge. We first train a default model on a roughly 88M-word corpus and test whether these preferences emerge in the dative. We then manipulate the input, balancing length and animacy related statistics in the training set to remove any order bias in the instances of the dative construction. We do this by creating a matching PO for every DO, and a matching DO for every PO. We show that dative preferences still persist—although they get weaker. So we train another model just like default except all datives are inserted in the reverse alternant, such that the direct evidence is now the opposite of what it normally would be. We find that this neutralizes the standard dative preferences. Then, we train models which do not have direct exposure to datives, and see that length preferences still emerge.
@@ 2.1 Corpus
Our training data is a subset of the 2023 BabyLM corpus with 100M words (Warstadt et al., 2023), which we choose for its tractable and human-scale size. We exclude the QED portion, since it includes special symbols and lacked proper sentence breaks, making dative detection difficult. This leaves our initial training set at around 88M words.
@@ 2.2 Defining and Detecting Datives
First, we obtained dependencies parses from spacy (Honnibal et al., 2020) for all utterances in the BabyLM dataset, on which the following definition of datives in the main paper is based. Then, we label an utterance as a DO dative if it contains a verb whose dependency contains a direct and an indirect object. For PO-datives, we consider utterances that have a verb with a dependency to a direct object (the theme), and dependencies to an indirect object (the recipient) of the following form: 1) dative or prep arc with to with a pobj to a pronoun or a noun phrase; 2) dative arc with for with a pobj to a pronoun or a noun phrase.
This definition is intentionally broad, with intent to cover all constructions of interest, but may include false positives, such as the PO-like burning it to the ground. Precision is less important than recall in controlled rearing experiments, since recall is key for controlling dative exposure in our models. More details, including an alternative (stricter) method for detecting datives and a demonstration that our results do not depend heavily on the precise operationalization of datives, can be found in App. F.
@@
Following Misra & Mahowald (2024), we artificially “pollute” our data with counterbalanced data intended to offset the effect of the undetected datives. These undetected datives could, in principle, provide direct evidence for human-like dative preferences. To counteract them, we insert datives, in roughly equal number, that violate the standard preferences. Specifically, for each model where we manipulate just the dative exposure, we insert 2250 counterfactual datives (swapping the structure of attested datives) to minimize the effects of the estimated 2250 false negative datives. We estimate that the undetected datives are twice as likely to be a DO than a PO, so 1500 counterfactuals are inserted in the DO form (when attested was PO) and 750 in PO form (when attested was DO). For each model, we will discuss if counterfactual insertion is done as they are introduced. See Table 7 in the Appendix for a breakdown of the number of controlled dative exposures, estimates of unaccounted-for datives, and insertions of counterfactuals by model.
@@
1a balanced I gave the dog a bone
I gave the dog a bone
I gave a bone to the dog
1b no-datives
I gave the dog a bone
He eats the green melon from the shop with a fork
I gave the dog a bone
He eats the green melon from the shop with a fork
no-2postverbal
I gave the dog a bone
He eats the green melon from the shop with a fork
I gave the dog a bone
He eats the green melon from the shop with a fork
1c swapped-datives I gave the dog a bone
I gave the dog a bone
I gave a bone to the dog
@@ 2.3 Model Architecture
We use the OPT-125M architecture (Zhang et al., 2022). See App. A for specific training details. Each of our LMs is trained using the transformers library (Wolf et al., 2020). We train each LM with three random seeds to account for randomness in weight initialization. For models with exposure to datives, we ensure that the total number of datives (including estimated unaccounted-for datives and counterfactual insertions) is equal across models, so that we can fairly compare dative preferences across models.
@@ 2.5 A Test Set for Testing Preferences by Feature
For future regression analyses, we compute a length difference score and an animacy difference score for each entry in the test set. Length difference is defined as the log difference in length between the recipient and theme. Because animacy is binary-coded, their difference scores take on values of $-1$, 0, or 1. An animacy difference of 0 means that either the theme and recipient are both animate or both inanimate, a difference of 1 means that the recipient is animate and theme is inanimate, and a difference of $-1$ means that the recipient is inanimate and the theme is animate.
The noun phrases in the attested datives are approximately in the natural distribution of features in DO and PO datives since they are randomly sampled. The average log difference in length between the recipient and theme is $-0.886$ in DOs and $0.199$ in POs, consistent with the easy-first preference. Animates are strongly aligned with the recipient position in either alternant (recipients were animate in 95.1% of DOs, and 71% of POs), consistent with the fact that inanimates are bad recipients (Beavers, 2011).
@@ 3 Preliminary Experiment: Does a BabyLM-trained LM learn human-like dative preferences?
While many dative verbs alternate freely between the DO and the PO, human alternation behavior for these verbs is graded (Bresnan & Nikitina, 2009). As a preliminary experiment, we test the extent to which our BabyLM-trained LMs capture human-like dative preferences in general. Specifically, we train an LM on our subset of the BabyLM corpus (see § 2.3), which we will refer to as default. The training set is filtered to consist of 66,822 datives in both DO and PO forms (and possibly undetected datives). The model should have roughly equal exposure to each alternant.
@@
For testing, we use the DAIS dataset from Hawkins et al. (2020), which contains 5000 different pairs of sentences in the DO and PO forms, spanning 200 different verbs. In addition, it also contains behavioral results from LMs such as GPT-2small and other models. We only select the sentences for which the dative verb appears in a dative construction in the training set, giving us 155 unique verbs and 3672 pairs of sentences. For each instance, the dataset also has a measure of the human judgments of preference for DO over PO sentences, analogous to our main measure of DO preference (see § 2.4). To compare our model’s preferences to those from humans, we followed Hawkins et al. and compared the average Human DO preference for the 155 target verbs to the average (z-scored) LM DO preference by computing their Pearson correlation.
Human-like Preferences Emerge
@@
We compute the correlations with human judgments obtained from our model and, for reference, we also re-compute the correlation obtained for GPT-2small (Radford et al., 2019) for the subset of the verbs we included in our analysis.  As seen in Fig. 2, our model (with seed selected by best evaluation loss) obtains a correlation of $r=0.50$ with the human ratings on DAIS. Not only is this positive, it is not far off from the correlation obtained by GPT-2small ($r=0.59$), even though our model was trained on a corpus that is several orders of magnitude smaller than that of GPT-2small.
@@
To measure the effect of length on dative preference, we compute the log-difference in length between the theme and recipient (positive value means recipient is shorter), and examine its correlation with the DO preference. As expected, we observe a negative correlation between the length difference and the DO preference: $r=-0.43$ for the default model with the best loss—i.e, the model prefers DO more when the recipient is shorter than the theme. We also found an animacy-first trend, such that DO preference was highest when the recipient was more animate than the theme.
[…]
@@ 4 Experiment 1: Balancing, Swapping, and Removing Datives
Having established that our small models show human-like dative preferences and, in particular, show a strong preference for length, a key question emerges: are these preferences a result of memorization from similar dative sentences, or do they emerge from a more general property of the input? To address this, we run a series of controlled rearing experiments, in which we systematically modify the evidence that the model receives and examine the persistence of the length and animacy effects.
In Experiment 1a, we manipulate the dative sentences in training to be perfectly balanced, effectively removing all length and animacy effects from dative constructions exposed to the model. In Experiment 1b, we use the same training set as for default, except all datives are in swapped-order (i.e. recipient and theme are “out-of-order”). This means that the model will receive the opposite direct evidence that we expect in the default case. In Experiment 1c, we then entirely remove dative constructions, and further all constructions with two postverbal arguments, so that the models do not receive any direct evidence. We also artificially pollute (see § 2.2) the training set to approximately neutralize the effect of undetected datives in Experiments 1a and 1c (skipping it in Experiment 1b since the vast majority of datives in its training set are in reversed order anyway).
Broadly, we assess the significance of our manipulations in two ways. The first is that we fit a mixed effect regression separately for each model, predicting DO preference based on length and animacy (as described above). This tells us whether, in each model, we observe length and animacy effects significantly different from 0 (Table 2 for all experiments). The second broad analysis is a global mixed effect model predicting DO preference across all experiments, based on fixed effects of length and animacy and the interact of each with model, treating default as baseline, as described in App. C. The interaction of each term (length and animacy) with model tells us whether our manipulations result in length/animacy effects significantly different from the default model (Table 4).
Figure 3: Correlation between DO preference and length difference (top) and animacy difference (bottom) by model on instances in the test set ($N=2000$). A negative correlation indicates short-first preference, while greater DO preference for recipients being animate indicates animate-first preference. The seeds with best evaluation loss are shown for length, whereas all three seeds are shown for animacy.

### v3

@@ Abstract
Language models (LMs) tend to show human-like preferences on a number of syntactic phenomena, but the extent to which these are attributable to direct exposure to the phenomena or more general properties of language is unclear. We explore this with the English dative alternation (DO: gave Y the X vs. PO: gave the X to Y), using a controlled rearing paradigm wherein we iteratively train small LMs on systematically manipulated input. We focus on two properties that affect the choice of alternant: length and animacy. Both properties are directly present in datives but also reflect more global tendencies for shorter elements to precede longer ones and animates to precede inanimates. First, by manipulating and ablating datives for these biases in the input, we show that direct evidence of length and animacy matters, but easy-first preferences persist even without such evidence. Then, using LMs trained on systematically perturbed datasets to manipulate global length effects (re-linearizing sentences globally while preserving dependency structure), we find that dative preferences can emerge from indirect evidence. We conclude that LMs’ emergent syntactic preferences come from a mix of direct and indirect sources.
@@ 1 Introduction
Yet most people would prefer the latter. The reason is that users of English typically show a “short-first” preference in the dative alternation. That is, if the theme is longer than the recipient they prefer the DO. If the theme is shorter, the PO. See an illustration of this in Fig.˜1. Besides length, there are a host of other factors that influence the choice of alternation used, such as animacy and pronominality (Bresnan et al., 2007; de Marneffe et al., 2012).
@@
These preferences in the dative alternation are learned well by even relatively small LMs (Hawkins et al., 2020; Misra & Kim, 2024). But what has to be true of the training data for these preferences to emerge? Recent work has considered the role of indirect evidence in language learning, asking whether LMs learn syntactic phenomena by memorizing only the most similar instances or whether they rely on structurally similar instances in order to generalize. Evidence so far has been mixed, with some work arguing for generalization from indirect evidence. Misra & Mahowald (2024) find that a somewhat rare syntactic structure is learned, in part, from indirect evidence from more common related structures. Patil et al. (2024) find that grammatical generalizations are possible even when never encountered. Oba et al. (2024), though, find that indirect evidence is less important than direct evidence. Houghton et al. (2025) find that binomial preferences (e.g., “peanut butter and jelly” vs. “jelly and peanut butter”) in models largely depend on memorization, with little evidence of generalization to human-like preferences for novel binomial pairs. Jumelet et al. (2024) find that language models can make some generalization about adjective order preference (e.g. “big black boxes” vs “black big boxes”) on adjective pairs not seen in training.
@@
We posit that the dative alternation is an elegant testing ground for exploring these questions, since it is linguistically well-studied and shown to be sensitive to a variety of factors that are not only directly observable in extant dative data, but which also depend on more general constraints in the language. One possibility (direct evidence) is that LMs prefer dative constructions that are highly similar to those observed in training. That is, if She gave it to the boy who signed up for class and was excited is attested in the training but She gave the boy who signed up for class and was excited it is not, then LMs might assign higher probability to strings that are very similar to the former. A second possibility (indirect evidence) is that the observed dative preferences are emergent from more general features of the input data. Specifically, we hypothesize that the LMs’ preference for “short-first” and “animate-first” dative utterances might follow not from memorizing instances of such datives but from those preferences existing more generally in the data (Behaghel, 1932; MacDonald, 2013) – e.g., Tur et al. (2025) find contemporary LMs to follow the general “short-before-long” ordering across a variety of syntactic structures.
@@
To adjudicate between these possibilities, we use a controlled rearing paradigm (Jumelet et al., 2021; Frank, 2023; Misra & Mahowald, 2024; Feng et al., 2024; Patil et al., 2024; Leong & Linzen, 2024; Kallini et al., 2024; Xu et al., 2025), whereby we train small models on carefully controlled input data and then test whether the predicted preferences emerge. We first train a default model on a roughly 88M-word corpus and test whether these preferences emerge in the dative. We then manipulate the input, balancing length and animacy-related statistics in the training set to remove any order bias in the instances of the dative construction. We do this by creating a matching PO for every DO, and a matching DO for every PO. We show that dative preferences still persist—although they get weaker. So we train another model just like default except all datives are inserted in the reverse alternant, such that the direct evidence is now the opposite of what it normally would be. We find that this neutralizes the standard dative preferences. Then, we train models which do not have direct exposure to datives, and see that length preferences still emerge.
@@ 2.1 Corpus
Our training data is a subset of the 2023 BabyLM corpus with 100M words (Warstadt et al., 2023), which we choose for its tractable and human-scale size. We exclude the QED portion, since it includes special symbols and lacks proper sentence breaks, making dative detection difficult. This leaves our initial training set at around 88M words.
@@ 2.2 Defining and Detecting Datives
{dependency}{deptext}
[column sep=1.5em] verb & recipient & theme
[…]
{dependency}{deptext}
[column sep=1.5em] verb & theme & to/for & recipient
[…]
Figure 2: spaCy parses of possible DO datives (top) and PO datives (bottom).
@@
First, we obtained dependencies parses from spaCy (Honnibal et al., 2020) for all utterances in the BabyLM dataset, on which the following definition of datives in the main paper is based. Then, we label an utterance as a DO dative if it contains a verb whose dependency contains a direct and an indirect object. For PO datives, we consider utterances that have a verb with a dependency to a direct object (the theme), and dependencies to an indirect object (the recipient) of the following form: 1) a to prepositional phrase under a dative or prep dependency; 2) a for prepositional phrase under a dative dependency. See Fig.˜2 for an illustration.
This definition is intentionally broad, with intent to cover all constructions of interest, but may include false positives, such as the PO-like burning it to the ground. Precision is less important than recall in controlled rearing experiments, since recall is key for controlling dative exposure in our models. More details, including an alternative (stricter) method for detecting datives and a demonstration that our results do not depend heavily on the precise operationalization of datives, can be found in App.˜F.
@@
Following Misra & Mahowald (2024), we artificially “pollute” our data with counterbalanced data intended to offset the effect of the undetected datives. These undetected datives could, in principle, provide direct evidence for human-like dative preferences. To counteract them, we insert datives, in roughly equal number, that violate the standard preferences. Specifically, for each model where we manipulate just the dative exposure, we insert 2250 counterfactual datives (swapping the structure of attested datives) to minimize the effects of the estimated 2250 false negative datives. We estimate that the undetected datives are twice as likely to be a DO than a PO, so 1500 counterfactuals are inserted in the DO form (when attested was PO) and 750 in PO form (when attested was DO). For each model, we will discuss if counterfactual insertion is done as they are introduced. See Table˜7 in the Appendix for a breakdown of the number of controlled dative exposures, estimates of unaccounted-for datives, and insertions of counterfactuals by model.
@@
1a balanced I gave the dog a bone I gave the dog a bone I gave a bone to the dog
1b swapped-datives I gave the dog a bone I gave the dog a bone I gave a bone to the dog
1c no-datives I gave the dog a bone He eats the green melon from the shop with a fork I gave the dog a bone He eats the green melon from the shop with a fork
no-2postverbal I gave the dog a bone He eats the green melon from the shop with a fork I gave the dog a bone He eats the green melon from the shop with a fork
@@ 2.3 Model Architecture
We use the OPT-125M architecture (Zhang et al., 2022). See App.˜A for specific training details. Each of our LMs is trained using the transformers library (Wolf et al., 2020). We train each LM with three random seeds to account for randomness in weight initialization. For models with exposure to datives, we ensure that the total number of datives (including estimated unaccounted-for datives and counterfactual insertions) is equal across models, so that we can fairly compare dative preferences across models.
@@ 2.5 A Test Set for Testing Preferences by Feature
For future regression analyses, we compute a length difference score and an animacy difference score for each entry in the test set. Length difference is defined as the log difference in length between the recipient and theme. Because animacy is binary-coded, their difference scores take on values of 1-1, 0, or 1. An animacy difference of 0 means that either the theme and recipient are both animate or both inanimate, a difference of 1 means that the recipient is animate and theme is inanimate, and a difference of 1-1 means that the recipient is inanimate and the theme is animate.
The noun phrases in the attested datives are approximately in the natural distribution of features in DO and PO datives since they are randomly sampled. The average log difference in length between the recipient and theme is 0.886-0.886 in DOs and $0.199$ in POs, consistent with the easy-first preference. Animates are strongly aligned with the recipient position in either alternant (recipients were animate in 95.1% of DOs, and 71% of POs), consistent with the fact that inanimates are bad recipients (Beavers, 2011).
@@ 3 Preliminary Experiment: Does a BabyLM-trained LM learn human-like dative preferences?
While many dative verbs alternate freely between the DO and the PO, human alternation behavior for these verbs is graded (Bresnan & Nikitina, 2009). As a preliminary experiment, we test the extent to which our BabyLM-trained LMs capture human-like dative preferences in general. Specifically, we train an LM on our subset of the BabyLM corpus (see §˜2.1), which we will refer to as default. The training set is filtered to consist of 66,822 datives in both DO and PO forms (and possibly undetected datives). The model should have roughly equal exposure to each alternant.
@@
For testing, we use the DAIS dataset from Hawkins et al. (2020), which contains 5000 different pairs of sentences in the DO and PO forms, spanning 200 different verbs. In addition, it also contains behavioral results from LMs such as GPT-2small and other models. We only select the sentences for which the dative verb appears in a dative construction in the training set, giving us 155 unique verbs and 3672 pairs of sentences. For each instance, the dataset also has a measure of the human judgments of preference for DO over PO sentences, analogous to our main measure of DO preference (see §˜2.4). To compare our model’s preferences to those from humans, we followed Hawkins et al. and compared the average Human DO preference for the 155 target verbs to the average (z-scored) LM DO preference by computing their Pearson correlation.
Clear Preferences Emerge
@@
We compute the correlations with human judgments obtained from our model and, for reference, we also re-compute the correlation obtained for GPT-2small (Radford et al., 2019) for the subset of the verbs we included in our analysis.  As seen in Fig.˜3, our model (with seed selected by best evaluation loss) obtains a correlation of $r=0.50$ with the human ratings on DAIS. Not only is this positive, it is not far off from the correlation obtained by GPT-2small ($r=0.59$), even though our model was trained on a corpus that is several orders of magnitude smaller than that of GPT-2small.
@@
To measure the effect of length on dative preference, we compute the log-difference in length between the theme and recipient (positive value means recipient is shorter), and examine its correlation with the DO preference. As expected, we observe a negative correlation between the length difference and the DO preference: r=0.43r=-0.43 for the default model with the best loss—i.e, the model prefers DO more when the recipient is shorter than the theme. We also found an animacy-first trend, such that DO preference was highest when the recipient was more animate than the theme.
We fit a mixed-effect model predicting DO preference based on fixed effects of length difference and animacy difference. We include random intercepts for verb lemma and seed with maximal random effect slope structures, excluding correlations to help convergence, following Barr et al. (2013). We compute p-values using nested model comparisons. For the default model, we found significant effects of length ($\hat{\beta}=-.174$, $p<.0001$) but not for animacy ($\hat{\beta}=.065$, $p=.11$). On further inspection, we found that the observed animate-first effects seem to be driven by specific verbs—e.g., leave it to me over leave me it, but write me a letter over write a letter to me. We include a discussion about verb-specific effects in App.˜E. We show length and animacy effects for default (and subsequent models) in Fig.˜4.
Figure 4: Correlation between DO preference and length difference (top) and animacy difference (bottom) by model on instances in the test set ($N=2000$). A negative correlation indicates short-first preference, while greater DO preference for recipients being animate indicates animate-first preference. The seeds with best evaluation loss are shown for length, whereas all three seeds are shown for animacy. Models shown here are for Experiment 1.
@@ 4 Experiment 1: Balancing, Swapping, and Removing Datives
Having established that our small models show clear dative preferences and, in particular, show a strong preference for length, a key question emerges: are these preferences a result of memorization from similar dative sentences, or do they emerge from a more general property of the input? To address this, we run a series of controlled rearing experiments, in which we systematically modify the evidence that the model receives and examine the persistence of the length and animacy effects.
In Experiment 1a, we manipulate the dative sentences in training to be perfectly balanced, effectively removing all length and animacy effects from dative constructions exposed to the model. In Experiment 1b, we use the same training set as for default, except all datives are in swapped-order (i.e. recipient and theme are “out-of-order”). This means that the model will receive the opposite direct evidence that we expect in the default case. In Experiment 1c, we then entirely remove dative constructions, as well as all constructions with two postverbal arguments, so that the models do not receive any direct evidence. We also artificially pollute (see §˜2.2) the training set to approximately neutralize the effect of undetected datives in Experiments 1a and 1c (skipping it in Experiment 1b since the vast majority of datives in its training set are in reversed order anyway).
Broadly, we assess the significance of our manipulations in two ways. The first is that we fit a mixed effect regression separately for each model, predicting DO preference based on length and animacy (as described above). This tells us whether, in each model, we observe length and animacy effects significantly different from 0 (Table˜2 for all experiments). The second broad analysis is a global mixed effect model predicting DO preference across all experiments, based on fixed effects of length and animacy and the interact of each with model, treating default as baseline, as described in App.˜C. The interaction of each term (length and animacy) with model tells us whether our manipulations result in length/animacy effects significantly different from the default model (Table˜4).